# SAD-Flower: Flow Matching for Safe, Admissible, and Dynamically Consistent Planning

### v1

@@ 5.2 Control Design using Control Lyapunov and Barrier Functions
where $\dot{V}(\bm{\tau}_{t})=\nabla^{T}V(\bm{\tau}_{t})\left(\bm{v}_{t}^{\bm{\theta}}(\bm{\tau}_{t})+\bm{u}_{t}\right)$ (CLF) and $\varphi(t)$ is a scheduling function that we will design later. Note that the computation of $\nabla^{T}V(\bm{\tau}_{t})$ requires knowledge of $\bm{f}$, but a model of it can be learned from the trajectory data if the true dynamical system in eq. (1) is unknown.
@@ 6.1 Experiment setting
We evaluate all methods in 4 different benchmark tasks: Maze2d, Hopper, and Walker2d from the D4RL problem set (maze_env), and Kuka Block-Stacking (janner2022planning).  [^3] Details of the experimental setting are provided in the Appendix. D
@@
6.2 Constrained-Planning Performance Across Benchmarks
As shown in Table 1, our method consistently satisfies safety and admissibility constraints while matching the planning performance of other methods across tasks. Although perfect dynamical consistency is not reached, the remaining violations are minor and mainly due to numerical integration of eq. (4). These results demonstrate the effectiveness of our control-theoretic mechanism, which guarantees constraint satisfaction at test time.
6.3 Why Does SAD-Flower Work Effectively?
QP vs. Non-convex Optimization. To assess the benefits of our QP-based formulation, we compare SAD-Flower with DPCC (romer2024diffusion), a recent method that also integrates optimal control into generative planning to enforce safety, admissibility, and dynamic consistency. DPCC applies control from the earliest stages and relies on full-system non-convex optimization at every sampling step. As shown in Table 2 for Hopper (Med-Expert), both methods achieve comparable levels of admissibility and dynamic consistency, but DPCC incurs much higher computation per sampling step and lower rewards.
@@ 6.3 Why Does SAD-Flower Work Effectively?
This highlights the efficiency of SAD-Flower gained without any loss of constraint satisfaction by enabling the usage of QP solvers compared to general non-convex optimization methods. Note though that the enforcement of constraints does still not come for free with QPs since the integration steps of the uncontrolled ODE merely take $0.01$s on average.
@@ 7 Conclusion
Datasets. We evaluate our method on a variety of public datasets across locomotion, navigation, and manipulation tasks. Full descriptions of the datasets can be found in Section 6 and Appendix D.
Compute. All experiments were conducted using NVIDIA Tesla P100 GPUs. Additional information on hardware and computational resources is provided in Appendix G.
@@ B.1 Feasibility of CBF constraints
Since the constraints defined in eqs. (CBF-s) and (CBF-a) concern independent variables, individual feasibility of each constraint implies joint feasibility of all of them. Due to the assume differentiability of $h_{k}^{s,a}$ and the definition of SDFs implying $\nabla h_{k}^{s,a}(\bm{\tau})||=\tfrac{1}{2}$, $\bm{u}$ can always be chosen such $\nabla^{T}h_{k}^{s,a}(\bm{\tau})\bm{u}$ takes an arbitrary value. Thus, the constraints defined in eqs. (CBF-s) and (CBF-a) are always feasible. ∎
@@ Appendix C Additional Experiment Details and Results
Appendix D Environment details
@@
Appendix E Constraint Settings in Each Tasks
@@
Appendix F Training detail
@@
Appendix G Computational resources
@@ 4 AMD EPYC 7542 NVIDIA Tesla P100 16 GB
Appendix H The Use of Large Language Models (LLMs)

### v2

@@ 5.2 Control Design using Control Lyapunov and Barrier Functions
where $\dot{V}(\bm{\tau}_{t})=\nabla^{T}V(\bm{\tau}_{t})\left(\bm{v}_{t}^{\bm{\theta}}(\bm{\tau}_{t})+\bm{u}_{t}\right)$ (CLF) and $\varphi(t)$ is a scheduling function that we will design later. Computing $\nabla^{T}V(\bm{\tau}_{t})$ requires access to $\bm{f}$. In some robotic domains, this information is available through high-fidelity simulators (gaz2019dynamic; howell2022dojo; acosta2022validating). If there exists no accurate physics-based simulator, e.g., for contact-rich manipulation tasks, a dynamics model can also be learned from the trajectory data. The learning errors directly correlate to the magnitude of dynamic consistency violations (curi2020efficient), such that practical consistency can still be achieved given a sufficiently precise learned model, and the (eq. (SC)), (eq. (AC)) can still be guaranteed, where is discussed in Section B.4.
@@ 6.1 Experiment setting
We evaluate all methods in 4 different benchmark tasks: Maze2d, Hopper, and Walker2d from the D4RL problem set (maze_env), and Kuka Block-Stacking (janner2022planning).  [^3] Details of the experimental setting are provided in the Appendix. E
@@
6.2 Constrained-Planning Performance Across Benchmarks
As shown in Table 1, our method consistently satisfies safety and admissibility constraints while matching the planning performance of other methods across tasks. Although perfect dynamical consistency is not reached, the remaining violations are minor and mainly due to numerical integration of eq. (4). These results demonstrate the effectiveness of our control-theoretic mechanism, which guarantees constraint satisfaction at test time. SAD-Flower achieves this performance at the cost of merely a small computation time increase compared to existing methods (Appendix, Table 15).
6.3 Why Does SAD-Flower Work Effectively?
@@
QP vs. Non-Convex Optimization. To assess the benefits of our QP-based formulation, we compare SAD-Flower with DPCC (romer2024diffusion), a recent method that also integrates optimal control into generative planning to enforce safety, admissibility, and dynamic consistency. DPCC applies control from the earliest stages and relies on full-system non-convex optimization at every sampling step. As shown in Table 2 for Hopper (Med-Expert), both methods achieve comparable levels of admissibility and dynamic consistency, but DPCC incurs much higher computation per sampling step and lower rewards. This highlights the efficiency of SAD-Flower gained without any loss of constraint satisfaction by enabling the usage of QP solvers compared to general non-convex optimization methods. Note though that the enforcement of constraints does still not come for free with QPs since the integration steps of the uncontrolled ODE merely take $0.01$s on average.
@@ 6.4 How reliable is SAD-Flower?
Table 3: Performance of SAD-Flower and FM in a dexterous grasping scenario.
Experiment Metric FM Ours
Adroit-Hand (Relocate) safety $15\pm 0.21 0.0$$00\textbf$$\bm{\pm}\textbf{0.00}$
admissib. $62\pm 0.19 0.0$$00\textbf$$\bm{\pm}\textbf{0.00}$
[…]
reward $\textbf{1.07}\pm\textbf{0.08}$ $1.05\pm 0.05$
Scalability in High-Dimensional Tasks. To evaluate the scalability of our approach in complex, contact-rich manipulation settings, we additionally consider the D4RL Adroit Relocate task, which involves a 39-dimensional state space and a 30-dimensional action space. This dexterous grasping environment poses significant challenges due to the need for precise contact interactions and high-dimensional control. We define the state constraint via the arm’s position to avoid collisions and the action constraint through actuator limits. As shown in Table 3, standard flow matching achieves moderate task success but suffers from frequent violations of safety and admissibility constraints due to unconstrained sampling. In contrast, SAD-Flower consistently satisfies both state and action constraints, demonstrating its ability to scale to high-dimensional robotic systems while maintaining constraint enforcement. This result underscores the robustness and generality of our control-augmented framework in realistic manipulation tasks.
Sensitivity Analysis of Constraint-Enforcement Strength. We conduct an ablation study on the LargeMaze task to assess the sensitivity of SAD-Flower to the hyperparameter $c$, which modulates the strength of constraint enforcement. While the main experiments use $c=0.5$, we vary it from $1.0$ to $0.2$. As shown in Table 4, SAD-Flower maintains perfect safety and admissibility across all tested values. Violations of dynamic consistency remain negligible and stable, while task performance varies moderately. These results demonstrate that SAD-Flower achieves constraint satisfaction without requiring precise hyperparameter tuning, highlighting its practical reliability.
Robustness under Imperfect Dynamics Models. To evaluate the robustness of our method under model approximation errors, we simulate degraded dynamics by training the forward model on progressively smaller subsets of the Hopper-Medium dataset. As shown in Table 5, SAD-Flower maintains full constraint satisfaction even with only $10\%$ of the original data, and dynamic consistency remains stable across these settings. Degradation becomes apparent only when the dataset is reduced to $0.01\%$, at which point safety violations emerge. These results demonstrate that SAD-Flower is tolerant to moderately inaccurate dynamics models and that training a sufficiently accurate model is feasible in practice.
Table 4: Sensitivity of SAD-Flower to the hyperparameter $c$.
Experiment Metric 1.0 0.8 0.6 0.4 0.2
Maze2d (Large) safety $\textbf{0.00}\bm{\pm}\textbf{0.00}$ $\textbf{0.00}\bm{\pm}\textbf{0.00}$ $\textbf{0.00}\bm{\pm}\textbf{0.00}$ $\textbf{0.00}\bm{\pm}\textbf{0.00}$ $\textbf{0.00}\bm{\pm}\textbf{0.00}$
admissib. $\textbf{0.00}\bm{\pm}\textbf{0.00}$ $\textbf{0.00}\bm{\pm}\textbf{0.00}$ $\textbf{0.00}\bm{\pm}\textbf{0.00}$ $\textbf{0.00}\bm{\pm}\textbf{0.00}$ $\textbf{0.00}\bm{\pm}\textbf{0.00}$
dyn. consist. $\textbf{0.01}\bm{\pm}\textbf{0.01}$ $\textbf{0.01}\bm{\pm}\textbf{0.01}$ $\textbf{0.01}\bm{\pm}\textbf{0.01}$ $\textbf{0.01}\bm{\pm}\textbf{0.01}$ $0.02\bm{\pm}0.01$
reward $1.38\pm 0.59$ $\textbf{1.54}\bm{\pm}\textbf{0.59}$ $1.46\pm 0.52$ $1.47\pm 0.53$ $1.29\bm{\pm}0.73$
Table 5: Effect of training dataset size on dynamic consistency and constraint satisfaction.
Experiment Metric 90% 70% 50% 30% 10% 0.01%
Hopper (Medium) safety $\pm 0.00 0.00±0.00$$\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $00\pm 0.00 0.0$$02\pm 0.09$
admissib. $\pm 0.00 0.00±0.00$$\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$
dyn. consist. $\pm 0.02 0.01±0.01$$\pm 0.01$ $01\pm 0.01 0.0$$1\pm 0.02 0.0$$1\pm 0.01 0.0$$04\pm 0.02$
reward $5\pm 0.01 0.33±0$$3\pm 0.04$ $0.35\pm 0.07$ $0.39\pm 0.07$ $0.38\pm 0.03$ $0.36\pm 0.03$
@@ 7 Conclusion
Datasets. We evaluate our method on a variety of public datasets across locomotion, navigation, and manipulation tasks. Full descriptions of the datasets can be found in Section 6 and Appendix E.
Compute. All experiments were conducted using NVIDIA Tesla P100 GPUs. Additional information on hardware and computational resources is provided in Appendix H.
@@ B.1 Feasibility of CBF constraints
Since the constraints defined in eqs. (CBF-s) and (CBF-a) concern independent variables, individual feasibility of each constraint implies joint feasibility of all of them. Due to the assumed differentiability of $h_{k}^{s,a}$ and the definition of SDFs implying $\nabla h_{k}^{s,a}(\bm{\tau})||=1$, $\bm{u}$ can always be chosen such $\nabla^{T}h_{k}^{s,a}(\bm{\tau})\bm{u}$ takes an arbitrary value. Thus, the constraints defined in eqs. (CBF-s) and (CBF-a) are always feasible. ∎
@@ B.2 Feasibility of CLF Constraints
B.4 Learning Error of the Dynamics Model
When the true system dynamics $\bm{f}^{*}$ are unknown, SAD-Flower relies on a learned forward model $\bm{f}$ trained from data. Inevitably, this introduces approximation error, which affects the predicted next states during sampling. Let the worst-case learning error be bounded by
$\bm{f}^{*}(\bm{s}(k),\bm{a}(k))-\bm{f}(\bm{s}(k),\bm{a}(k))||\leq\zeta$ (10)
Let $\bm{s}^{*}(k)$ denote the state propagated by the true dynamics and $\bm{s}(k)$ the state propagated by the learned model:
$\bm{s}^{*}(k)=\bm{f}^{*}(\bm{s}^{*}(k-1),\bm{a}(k-1)),\qquad\bm{s}(k)=\bm{f}(\bm{s}(k-1),\bm{a}(k-1)).$
Their deviation satisfies
[…]
$\leq\xi,$ (11)
where $\xi=\zeta\sum_{i=0}^{k-1}L_{f}^{i}$ and $L_{f}$ is the Lipschitz constant of the learned forward model, which can be obtained with several techniques, e.g., automatic differentiation algorithms (virmaux2018lipschitz), convex optimization (fazlyab2019efficient).
Equation eq. (B.4) provides an explicit upper bound $\xi$ on the mismatch between the true and learned system trajectories. To ensure safety and admissibility for the true system despite this model error, we apply a robust CBF formulation (buch2021robust). The state-constraint CBF condition in Eq. CBF-s is modified as
[…]
where the $L_{h}$ and $L_{\nabla h}$ denote Lipschitz constants of $h^{s}_{k}(\bm{\tau}_{t})$ and its gradient.
With this modification, the robust control law for $t\geq T_{0}$ becomes
[…]
Under this controller (and $\bm{u}_{t}=0$ for $t<T_{0}$), the resulting trajectory $\bm{\tau}^{*}_{t}$ rendered by the true system satisfies the safety and admissibility constraints eq. (SC) and eq. (AC). Although dynamic consistency for the true system cannot be formally guaranteed without knowing $\bm{f}^{*}$, the robust CBF formulation ensures that state and action constraints remain satisfied even with imperfect learned dynamics.
@@ Appendix C Additional Experiment Details and Results
Appendix D Comparison with constraint-aware baselines
We further compare SAD-Flower against two additional constraint-aware baselines: a reject-sampling method, which offers a simple heuristic for constraint handling, and CoBL-Diffusion (mizuta2024cobl), which leverages control-theoretic rewards to encourage constraint satisfaction.
The reject-sampling approach first trains a flow matching model on the original dataset, which includes trajectories that violate test-time constraints, since those constraints are unseen during training. At inference, it samples a batch of trajectory candidates and discards those violating constraints, selecting from the remaining set the trajectory with the least violation. As shown in Table 8, this approach consistently fails to satisfy constraints across all tasks, including Maze2d, Hopper, Walker2d, and Kuka Block-Stacking. Both safety and admissibility violations remain non-zero. This confirms that post-hoc filtering alone is insufficient when the generative model has no knowledge of constraint structure—especially under novel test-time constraints. Even though flow matching can model multimodal behavior patterns, it cannot reliably sample within unseen constraint sets unless such constraints are explicitly incorporated during inference.
CoBL-Diffusion injects physical constraints into a diffusion model by shaping its denoising process with auxiliary rewards derived from Control Barrier Functions (CBFs) and Control Lyapunov Functions (CLFs). This allows it to softly bias generation toward constraint-adherent behaviors, but without any formal guarantees. We evaluate CoBL-Diffusion in the LargeMaze environment, using the same horizon and constraints as in other baselines. As shown in Table 9, the method fails to fully enforce safety and admissibility constraints. This is expected given several key limitations: (1) it predicts only actions, whereas SAD-Flower jointly generates states and actions, which allows for more precise control of trajectory behavior; (2) its CBFs target only state constraints, while ours apply to both state and action spaces; (3) it uses CLFs for general stability, not for ensuring formal dynamic consistency as we do; and (4) it lacks prescribed-time scheduling, making it difficult to guarantee constraint satisfaction at the final step. These design choices limit CoBL-Diffusion’s ability to robustly handle complex, temporally structured constraints. In contrast, SAD-Flower provides a framework to reliably enforce constraints even under novel test-time scenarios.
Table 8: Performance of SAD-Flower and reject-sample method across all tasks.
Experiment Metric Reject-sample Ours
Maze2d (Large) safety $16\pm 0.29 0.0$$00\textbf$$\bm{\pm}\textbf{0.00}$
admissib. $90\pm 0.01 0.0$$00\textbf$$\bm{\pm}\textbf{0.00}$
dyn. consist. $02\pm 0.01 0.0$$01\textbf$$\bm{\pm}\textbf{0.01}$
reward $\textbf{1.52}\pm\textbf{0.28}$ $1.42\pm 0.52$
Maze2d (Umaze) safety $4\pm 0.09 0.00±0$$0\textbf$$\bm{\pm}\textbf{0.00}$
admissib. $91\pm 0.01 0.0$$00\textbf$$\bm{\pm}\textbf{0.00}$
dyn. consist. $\textbf{0.01}\bm{\pm}\textbf{0.01}$ $\textbf{0.01}\bm{\pm}\textbf{0.01}$
reward $\textbf{2.91}\pm\textbf{0.65}$ $2.66\pm 0.88$
Hopper (Med-Expert) safety $14\pm 0.18 0.0$$00\textbf$$\bm{\pm}\textbf{0.00}$
admissib. $5\pm 0.02 0.00±0$$0\textbf$$\bm{\pm}\textbf{0.00}$
dyn. consist. $02\pm 0.01 0.0$$01\textbf$$\bm{\pm}\textbf{0.01}$
reward $0.49\pm 0.33$ $\textbf{0.93}\pm\textbf{0.23}$
Hopper (Medium) safety $3\pm 0.04 0.00±0$$0\textbf$$\bm{\pm}\textbf{0.00}$
admissib. $6\pm 0.03 0.00±0$$0\textbf$$\bm{\pm}\textbf{0.00}$
dyn. consist. $02\pm 0.02 0.0$$01\textbf$$\bm{\pm}\textbf{0.01}$
reward $0.24\pm 0.14$ $\textbf{0.34}\pm\textbf{0.03}$
Walker2d (Med-Expert) safety $8\pm 0.10 0.00±0$$0\textbf$$\bm{\pm}\textbf{0.00}$
admissib. $8\pm 0.08 0.00±0$$0\textbf$$\bm{\pm}\textbf{0.00}$
dyn. consist. $\textbf{0.04}\bm{\pm}\textbf{0.06}$ $\textbf{0.04}\bm{\pm}\textbf{0.04}$
reward $\textbf{0.97}\pm\textbf{0.22}$ $0.89\pm 0.32$
Walker2d (Medium) safety $18\pm 0.15 0.0$$00\textbf$$\bm{\pm}\textbf{0.00}$
admissib. $16\pm 0.10 0.0$$00\textbf$$\bm{\pm}\textbf{0.00}$
dyn. consist. $\textbf{0.07}\bm{\pm}\textbf{0.17}$ $\textbf{0.07}\bm{\pm}\textbf{0.15}$
reward $\textbf{0.68}\pm\textbf{0.19}$ $0.42\pm 0.23$
KUKA Block Stacking safety $1\pm 0.01 0.00±0$$0\textbf$$\bm{\pm}\textbf{0.00}$
reward $0.44\pm 0.67$ $\textbf{0.45}\pm\textbf{0.21}$
Table 9: Performance of SAD-Flower and CoBL-Diffusion in the navigation task.
Experiment Metric CoBL-Diffusion Ours
[…]
dyn. consist. $\bm{\pm}0.00$ $\pm0.01$
[…]
Appendix E Environment details
@@
Appendix F Constraint Settings in Each Tasks
@@
Appendix G Training detail
@@
Appendix H Computational resources
@@ 4 AMD EPYC 7542 NVIDIA Tesla P100 16 GB
Appendix I Computation Analysis
We report the detailed computational costs of SAD-Flower and all baseline methods across our benchmark tasks in Table 15. As expected, SAD-Flower incurs greater computational cost compared to unconstrained generative planners such as Diffuser and FM, due to the overhead introduced by enforcing multiple constraints during sampling. Nonetheless, in some domains—such as Hopper-Medium-Expert and Kuka Block-Stacking—SAD-Flower achieves lower runtime than SafeDiffuser, highlighting the efficiency gains enabled by our prescribed-time control formulation. Importantly, SAD-Flower is the only method that consistently satisfies all three constraint classes—state, action, and dynamic consistency—across all evaluated domains. No faster baseline provides this level of safety and reliability.
Table 15: Computation analysis of the proposed SAD-Flower and baselines across navigation, locomotion, and manipulation tasks.
Experiment Metric Diffuser Trunc CG FM S-Diffuser D-Diffuser Ours
Maze2d (Large) comp. time $01\pm 0.01 0.0$$2\pm 0.01 0.0$$4\pm 0.01 0.0$$6\pm 0.02 0.0$$6\pm 0.01 0.0$$01\pm 0.01$ $0.13\pm 0.01$
Maze2d (Umaze) comp. time $01\pm 0.01 0.0$$4\pm 0.01 0.0$$2\pm 0.01 0.0$$6\pm 0.02 0.0$$1\pm 0.01 0.0$$08\pm 0.02$
Hopper (Med-Expert) comp. time $05\pm 0.01 0.0$$6\pm 0.02 0.0$$6\pm 0.01 0.0$$\pm 0.01 0.11±0$$1\pm 0.02$ $02\pm 0.01 0.0$$06\pm 0.01$
Hopper (Medium) comp. time $05\pm 0.01 0.0$$6\pm 0.01 0.0$$6\pm 0.02 0.0$$1\pm 0.01 0.0$$9\pm 0.02 0.0$$02\pm 0.01$ $0.12\pm 0.02$
Walker2D (Med-Expert) comp. time $06\pm 0.02 0.0$$6\pm 0.01 0.0$$5\pm 0.01 0.0$$1\pm 0.01 0.0$$9\pm 0.02 0.0$$2\pm 0.01 0.0$$09\pm 0.01$
Walker2D (Medium) comp. time $0.04\pm 0.01$ $0.13\pm 0.03$ $21\pm 0.02 0.0$$01\pm 0.01$ $10\pm 0.02 0.0$$2\pm 0.01 0.0$$09\pm 0.01$
KUKA Block Stacking comp. time $0.70\pm 0.01$ $0.84\pm 0.01$ $0.86\pm 0.01$ $0.47\pm 0.06$ $78\pm 0.01 0.0$$04\pm 0.01$ $0.74\pm 0.08$
Appendix J Deployment on Real-World Robotic Platforms
Recent works have demonstrated the successful deployment of diffusion- and flow-based models on physical robotic systems (chi2023diffusion; sochopoulos2025fast), highlighting the feasibility of integrating such models into real-world robotic pipelines that involve perception, state estimation, and control. These results indicate that generative trajectory models can operate effectively outside of simulation.
SAD-Flower is designed as an open-loop planner that generates full trajectories prior to execution, without requiring real-time feedback control. This formulation avoids the need to meet strict control-frequency constraints, making it amenable to practical deployment. Although SAD-Flower solves a small quadratic program (QP) at each sampling step to enforce constraints, this overhead remains moderate and does not hinder offline trajectory generation. Moreover, recent advances in efficient generative planning—such as DiffuserLite (dong2024diffuserlite) and Habi (lu2025habitizing)—can be directly incorporated into our framework to further accelerate sampling, enabling tighter planning budgets without sacrificing constraint guarantees.
While our current implementation is focused on simulation, the modular structure of SAD-Flower makes it compatible with standard deployment pipelines. The forward model used for dynamics consistency can be learned directly from onboard sensor data, and constraint formulations (e.g., actuator limits, workspace boundaries) can be adapted to real-world specifications. As with all generative models, care must be taken to address sensor noise and execution uncertainty, but our control-theoretic formulation inherently provides robustness by enforcing constraints throughout the trajectory.
In summary, SAD-Flower introduces only modest additional computation to existing generative planners while providing formal guarantees for safety, admissibility, and dynamic consistency. Given the demonstrated deployability of similar frameworks and the compatibility of our method with efficient sampling and standard control pipelines, we believe SAD-Flower can be incorporated into real-world robotic systems with limited adaptation. Investigating such deployment remains an important direction for future work.
Appendix K The Use of Large Language Models (LLMs)